# Exact 3-D Channel Impulse Response Under Uniform Drift for Absorbing Spherical Receivers

### v2

@@ I Introduction
For three-dimensional molecular communication (MC) systems [3, 12, 5], the point-to-sphere absorbing channel has long served as a canonical reference model [7]. Its analytical tractability, however, relies critically on strong geometric symmetries that are immediately broken once drift is introduced.
In practical MC environments, drift is not a secondary effect but a fundamental transport mechanism, arising from background flow [15, 8], externally induced bias fields [11, 4], or spatial concentration gradients. Even when the drift field is spatially uniform, its presence breaks the radial symmetry that underlies classical no-drift formulations [17]. As a result, the first-hitting statistics depend not only on the arrival time but also on the angular location at which particles are absorbed.
Despite its physical relevance, an exact channel impulse response (CIR) for the point-to-sphere geometry under arbitrary drift direction has remained unavailable. Existing analytical results are restricted to zero-drift scenarios [17], aligned drift configurations [2], or reduced-dimensional approximations [15, 8], while more general settings are typically handled through Monte Carlo simulations or empirical fitting [7, 1]. This lack of an exact reference model has limited both physical interpretability and systematic channel analysis in drift-dominated regimes.
In this letter, we close this gap by deriving an exact analytical CIR for an absorbing spherical receiver [^1] In the molecular communication literature, a fully absorbing spherical receiver is widely used as a standard model for ligand–receptor binding or perfect capture at the receiver surface. in 3-D space under uniform drift with arbitrary direction. Rather than modifying existing no-drift expressions, we formulate the problem at the level of the joint first-hitting time and location statistics. By applying a measure-change argument based on the Girsanov theorem [14], the effect of drift is incorporated as an explicit reweighting of boundary hitting events, while preserving the analytical structure of the baseline no-drift solution.
@@
$d=1$ Lévy distribution [8, 15] Inverse Gaussian distribution [8, 15]
$d=3$ (Planar) Exact analytical CIR [10] Exact analytical CIR (arbitrary direction) [10]
@@ II System Model
We consider an unbounded 3-D fluid environment $\mathbb{R}^{3}$. The transmitter (Tx) is modeled as a point source located at $\mathbf{x}_{0}\in\mathbb{R}^{3}$, and the receiver (Rx) is a fully absorbing sphere of radius $r$ centered at the origin, denoted by $\mathcal{S}_{r}$, with $|\mathbf{x}_{0}|>r$. The system geometry is illustrated in Fig. 1.
Information particles (molecules) are released at time $t=0$. From a microscopic perspective, the trajectory of a particle $\mathbf{X}_{t}$ follows the Itô stochastic differential equation [14]
@@
While the proposed framework applies to general drift profiles, this letter focuses on the practically relevant case of constant uniform drift, $\mathbf{u}(t)\equiv\mathbf{v}\in\mathbb{R}^{3}$, forming an angle $\psi=\angle(\mathbf{v},\mathbf{x}_{0})$ with the Tx–Rx axis. The relative strength of advection and diffusion is characterized by the Péclet number $\mathrm{Pe}=|\mathbf{v}|r/D$. The first hitting time is defined as $T\triangleq\inf\{t>0:\ |\mathbf{X}_{t}|=r\},$ and the CIR is the probability density function of $T$, denoted by $f_{\mathrm{3D}}^{(\mathbf{v})}(t)$.
@@
Yilmaz et al. [17] 3-D No-Drift CIR (Marginal Time) Yin [18] Baseline 3-D Joint Time-Position The Drift-Factor Method Apply Girsanov Theorem (i.e., Radon-Nikodym Derivative) Proposed Theorem 1 Exact 3-D CIR with Arbitrary Drift Direction Insufficient (Lacks spatial info) Basis for Measure ChangeIntegration over Sphere
Figure 2: Derivation flow based on drift-factor method. We start from Yin’s joint time-position distribution [18] and apply the Girsanov theorem to derive the 3D CIR in analytical form.
@@ III-A Motivation: Why the Joint Distribution Is Necessary
Existing analytical results for spherical receivers without drift typically rely on the marginal first-hitting-time distribution [17]. However, under drift, advection introduces a strong directional bias: the contribution of each hitting event depends not only on when the particle hits the receiver, but also on where it hits the spherical boundary. Consequently, the marginal hitting-time distribution alone is insufficient to characterize the drifted CIR.
To retain this spatial information, we start from the joint probability density of the first hitting time $T$ and the hitting location $\mathbf{X}_{T}$ on the spherical boundary $\mathcal{S}_{r}$, denoted by $f_{\mathrm{3D}}^{(\mathbf{0})}(t,\mathbf{y})$ for $\mathbf{y}\in\mathcal{S}_{r}$. An explicit analytical expression for this joint density in the no-drift case is available from probability literature [18] and serves as the baseline of our derivation.
@@ III-C Baseline Joint Density for the No-Drift Case
For a Tx located outside the spherical Rx (i.e., $|\mathbf{x}_{0}|>r$), the joint density of the hitting time-location for a 3-D Brownian motion without drift admits the following series representation, adapted from [18, Theorem 3.2]:
$f_{\text{3D}}^{(\mathbf{0})}(t,\mathbf{y})$ $=\frac{1}{2\pi}\frac{r}{|\mathbf{x}_{0}|}\sum_{m=0}^{\infty}\Big(m+\tfrac{1}{2}\Big)P_{m}\!\big(\cos\angle(\mathbf{x}_{0},\mathbf{y})\big)$
$\times$ $\int_{0}^{\infty}\frac{\lambda\,Z_{m+\frac{1}{2}}(|\mathbf{x}_{0}|/r,\lambda r/\sigma)}{J_{m+\frac{1}{2}}^{2}(\lambda r/\sigma)+Y_{m+\frac{1}{2}}^{2}(\lambda r/\sigma)}\exp\!\left(-\tfrac{1}{2}\lambda^{2}t\right)\,d\lambda.$ (3)
@@ III-E Mode-Wise Surface Integration
To evaluate these integrals, we adapt the following identity from [18, Lemma 2.5] specialized to the 3-D setting.
@@
In $\mathbb{R}^{3}$, the CIR for a spherical absorbing receiver of radius $r$ under a uniform drift $\mathbf{v}$ is
@@
This exact representation (1) serves as the basis for the numerical evaluation and performance characterization presented in the next section.
@@
Figure 3: Validation of analytical CIR against Monte Carlo simulations. The analytical curves correspond to $N_{\text{tx}}f_{\text{3D}}^{(\mathbf{v})}(t)\,\Delta t$, i.e., the expected number of absorbed molecules within a fixed time bin $\Delta t=5\times 10^{-5}\,\mathrm{s}$. Top Row: Moderate drift speed ($|\mathbf{v}|=5\,\mu\text{m/s}$). Bottom Row: Higher drift speed ($|\mathbf{v}|=10\,\mu\text{m/s}$). Columns correspond to positive ($\psi=180^{\circ}$), transverse ($\psi=90^{\circ}$), and negative ($\psi=0^{\circ}$) drift directions.
@@ IV Numerical Results and Discussion
In this section, we validate the derived exact analytical CIR against particle-based Monte Carlo simulations and give a brief discussion on how the proposed formulation enables efficient characterization of two key channel metrics: pulse amplitude and pulse peak time. All numerical evaluations and simulations are performed using MATLAB (MathWorks cloud-based environment). To ensure consistency with established MCvD channel modeling literature [17], the default system parameters are chosen typically as follows: diffusion coefficient $D=80\,\mu\mathrm{m}^{2}/\mathrm{s}$, receiver radius $r=10\,\mu\mathrm{m}$, and initial distance $|\mathbf{x}_{0}|=20\,\mu\mathrm{m}$. We consider two representative drift scenarios: a moderate drift speed ($|\mathbf{v}|=5\,\mu\mathrm{m/s}$) and a higher drift speed ($|\mathbf{v}|=10\,\mu\mathrm{m/s}$).
For Monte Carlo validation shown in Fig. 3, the analytical CIR is discretized over a fixed time bin $\Delta t$ to enable a direct comparison with particle-based simulations. In Figs. 4 and 5, the peak time $t_{\mathrm{peak}}$ is obtained by maximizing the continuous-time analytical CIR, while the corresponding peak amplitude is reported in terms of the expected number of absorbed molecules per time bin, $N_{\mathrm{tx}}f^{(\mathbf{v})}_{\mathrm{3D}}(t_{\mathrm{peak}})\Delta t$, using the same $\Delta t$ as in Fig. 3 for consistent scaling.
@@ IV-B Convergence of the Analytical Series
Numerical evaluation of Theorem 1 requires truncation of the infinite series over the angular mode index $m$. From the structure of (1), the truncation order $M$ controls the angular resolution of the CIR through the Legendre modes $P_{m}(\cos\psi)$, while temporal smoothing is governed by the Gaussian factor $e^{-\lambda^{2}t/2}$ in the radial integral.
For moderate drift strengths, higher-order angular modes are increasingly suppressed by the modified Bessel term $I_{m+\frac{1}{2}}(|\mathbf{v}|r/\sigma^{2})$, and truncation at $M=30$ yields a relative error below $10^{-5}$ over most of the time domain for the parameters considered in this work. This accuracy is sufficient for Monte Carlo validation and for the extraction of peak metrics.
Convergence becomes slower only in the immediate vicinity of the first arrival (very small $t$), where temporal smoothing is weak and sharper angular features must be resolved, or under stronger drift conditions that induce pronounced angular anisotropy. In such cases, increasing the truncation order to $M\approx 50$ provides accurate resolution of the initial transient. For practical operating regimes, $M=30$ offers an effective accuracy–complexity trade-off.
@@ IV-C Characterization of Peak Metrics
Monte Carlo estimation of peak metrics is computationally expensive because extensive averaging is needed to suppress stochastic fluctuations. In contrast, the analytical CIR provides a smooth objective function, enabling efficient numerical maximization. Compared with particle-based Monte Carlo estimation, this approach reduces the computational cost by several orders of magnitude, as it avoids repeated stochastic sampling and averaging.
@@ V Conclusion
This letter establishes an exact analytical characterization of the 3-D CIR for MCvD systems with a fully absorbing spherical receiver under uniform drift of arbitrary direction. By formulating the problem at the level of joint first-hitting time and location statistics and applying a measure-change argument, we derive an exact series representation of the CIR within the drift–diffusion model, without relying on geometric approximations [2] or dimensional reductions [8]. The analytical CIR is validated against particle-based Monte Carlo simulations over representative drift magnitudes and orientations, showing excellent agreement. Beyond validation, the exact formulation enables efficient and noise-free extraction of key channel metrics via direct numerical maximization with respect to the CIR curve with substantially lower computational cost than simulation-based approaches. Numerical results further clarify how drift orientation and receiver geometry jointly shape the received signal, capturing enhancement, suppression, and geometric scaling effects in a unified manner. From a system-level perspective, the proposed CIR provides a precise and computationally tractable channel model that can be potentially integrated into molecular MIMO detection and receiver design (e.g., [6, 1]), completing a previously missing exact reference model for drifted point-to-sphere molecular channels [7].
Appendix A Methodological Consistency Check
This appendix provides a brief consistency check for the proposed measure-change framework. It is not used as a reduction or a derivation step for the 3-D spherical model in the main text.
In one dimension, an absorbing boundary implies a deterministic point hitting location on the hitting event. Let the initial distance be $\ell>0$ and $\sigma^{2}=2D$. The no-drift first-hitting-time density is $f_{\mathrm{1D}}^{(0)}(t)=\frac{\ell}{\sqrt{4\pi Dt^{3}}}\exp\!\left(-\frac{\ell^{2}}{4Dt}\right),~t>0$, see [15]. Under constant drift $v$ (towards the boundary), the Girsanov factor evaluated at the hitting time reduces to $\exp\!\left(\frac{v\ell}{2D}-\frac{v^{2}t}{4D}\right)$, corresponding to (2) with $\mathbf{v}\!\cdot\!(\mathbf{y}-\mathbf{x}_{0})=v\ell$ in one dimension. Multiplying this factor with the no-drift density $f_{\mathrm{1D}}^{(0)}(t)$ yields
$f_{\mathrm{1D}}^{(v)}(t)$ $=\frac{\ell}{\sqrt{4\pi Dt^{3}}}\exp\!\left(-\frac{\ell^{2}}{4Dt}+\frac{v\ell}{2D}-\frac{v^{2}t}{4D}\right)$
$=\frac{\ell}{\sqrt{4\pi Dt^{3}}}\exp\!\left(-\frac{(\ell-vt)^{2}}{4Dt}\right),$ (7)
where the exponent is obtained by completing the square. This recovers the standard inverse Gaussian first-passage-time law (cf. [8, 15]), which corresponds to the 1-D uniform-drift case as presented in Table I.

### v3

@@ I Introduction
The point-to-sphere absorbing channel [8] stands as the canonical reference model for three-dimensional (3-D) molecular communication (MC) systems [4, 14, 6]. In quiescent environments, this geometry allows for compact analytical characterizations of first-hitting statistics by exploiting its inherent radial symmetry [19]. However, practical MC scenarios are rarely static; drift arises naturally from background flow [17, 9], externally induced bias fields [13, 5], or concentration-driven transport.
The introduction of advection—even if spatially uniform—fundamentally breaks the radial symmetry that classical formulations rely upon. Unlike the no-drift case, advection creates a directional bias that couples the first-hitting time with the angular absorption location. This symmetry breaking introduces a level of analytical complexity that precludes the direct extension of radial-only models to flow-perturbed environments.
Consequently, despite its physical relevance, an exact analytical channel impulse response (CIR) for the point-to-sphere geometry under arbitrary drift direction has remained unavailable since the seminal work [19] of Yilmaz et al. in 2014. Existing results are restricted to zero-drift scenarios [19], specifically aligned drift configurations [2], or reduced-dimensional approximations [17, 9], while general settings are typically relegated to Monte Carlo simulations or empirical fitting [8, 1]. This lack of an exact reference model has limited both physical interpretability and systematic channel analysis in drift-dominated regimes.
In this letter, we close this gap by deriving an exact analytical CIR for an absorbing spherical receiver in 3-D space under uniform drift with arbitrary direction. Rather than modifying existing no-drift expressions, we formulate the problem at the level of the joint first-hitting time and location statistics. By applying a measure-change argument based on the Girsanov theorem [16], the effect of drift is incorporated as an explicit reweighting of boundary hitting events, while preserving the analytical structure of the baseline no-drift solution.
@@
$d=1$ Lévy distribution [9, 17] Inverse Gaussian (IG) distribution [9, 17]
$d=3$ (Planar) Exact analytical CIR [11, 12] Exact analytical CIR (arbitrary direction) [11, 12]
@@ II System Model
We consider an unbounded 3-D fluid environment $\mathbb{R}^{3}$. The transmitter (Tx) is modeled as a point source located at $\mathbf{x}_{0}\in\mathbb{R}^{3}$, and the receiver (Rx) is a fully-absorbing (FA) sphere of radius $r$ centered at the origin, denoted by $\mathcal{S}_{r}$, with $|\mathbf{x}_{0}|>r$. The system geometry is illustrated in Fig. 1.
Remark 1 (Practical relevance of the FA model).
The FA receiver model is a widely adopted transport-level abstraction within MC (see [8, Fig. 8]). Biochemically, it represents the limiting behavior ($\kappa_{f}\to\infty$) of high-affinity ligand–receptor binding, where molecules are effectively removed upon contact with the receiver surface; numerical studies indicate that even a modest receptor surface coverage (on the order of $1\%$) suffices to closely approximate the absorption statistics of an ideal FA sphere [3].
Information particles (molecules) are released at time $t=0$. From a microscopic perspective, the trajectory of a particle $\mathbf{X}_{t}$ follows the Itô stochastic differential equation [16],
@@
While the proposed measure-change framework applies to general drift profiles, this letter focuses on the practically relevant case of constant uniform drift, $\mathbf{u}(t)\equiv\mathbf{v}\in\mathbb{R}^{3}$, forming an angle $\psi=\angle(\mathbf{v},\mathbf{x}_{0})$ with the Tx–Rx axis. The relative strength of advection and diffusion is characterized by the Péclet number $\mathrm{Pe}=|\mathbf{v}|r/D$. The first hitting time is defined as $T\triangleq\inf\{t>0:\ |\mathbf{X}_{t}|=r\},$ and the CIR is the probability density function of $T$, denoted by $f_{\mathrm{3D}}^{(\mathbf{v})}(t)$.
@@
Yilmaz et al. [19] 3-D No-Drift CIR (Marginal Time) Yin [20] Baseline 3-D Joint Time–Position The Drift-Factor Method Apply Girsanov Theorem (i.e., Radon-Nikodym Derivative) Proposed Theorem 1 Exact 3-D CIR with Arbitrary Drift Direction Insufficient (Lacks spatial info) Basis for Measure ChangeIntegration over Sphere
Figure 2: Derivation flow based on drift-factor method. We start from Yin’s joint time-position distribution [20] and apply the Girsanov theorem to derive the 3-D CIR in analytical form.
@@ III-A Motivation: Why the Joint Distribution Is Necessary
Classical analytical results for spherical FA receivers without drift typically rely on the marginal first-hitting-time distribution [19], as radial symmetry renders the angular coordinate immaterial. Under drift, however, advection introduces directional bias that breaks this symmetry, so the channel response depends not only on when particles arrive but also on where they hit the spherical boundary. As a result, the marginal hitting-time distribution alone is insufficient to characterize the drifted CIR. To access this spatial information, we begin from the joint density of the first hitting time $T$ and the hitting location $\mathbf{X}_{T}$ on $\mathcal{S}_{r}$, denoted by $f_{\mathrm{3D}}^{(\mathbf{0})}(t,\mathbf{y})$, for which an explicit analytical expression in the no-drift case is available in the probability literature [20] and serves as the baseline of our derivation.
@@ III-C Baseline Joint Density for the No-Drift Case
For a Tx located outside the spherical Rx (i.e., $|\mathbf{x}_{0}|>r$), the joint density of the hitting time-location for a 3-D Brownian motion without drift admits the following series representation, adapted from [20, Theorem 3.2],
$f_{\mathrm{3D}}^{(\mathbf{0})}(t,\mathbf{y})$ $=\frac{1}{2\pi}\frac{r}{|\mathbf{x}_{0}|}\sum_{m=0}^{\infty}\Big(m+\tfrac{1}{2}\Big)P_{m}\!\big(\cos\angle(\mathbf{x}_{0},\mathbf{y})\big)$
$\times$ $\int_{0}^{\infty}\frac{\lambda\,Z_{m+\frac{1}{2}}(|\mathbf{x}_{0}|/r,\lambda r/\sigma)}{J_{m+\frac{1}{2}}^{2}(\lambda r/\sigma)+Y_{m+\frac{1}{2}}^{2}(\lambda r/\sigma)}\exp\!\left(-\tfrac{1}{2}\lambda^{2}t\right)\,d\lambda,$ (3)
@@ III-E Mode-Wise Surface Integration
To evaluate these integrals, we adapt the following identity from [20, Lemma 2.5] specialized to the 3-D setting.
@@
In $\mathbb{R}^{3}$, the CIR for a spherical absorbing receiver of radius $r$ under a uniform drift $\mathbf{v}$ is,
@@
Remark 2 (Numerical Convergence and Evaluation).
The series in (1) converges rapidly for $t>0$ because the modified Bessel function $I_{m+1/2}(z)$ decays at a factorial rate (specifically, $I_{\nu}(z)\sim\frac{(z/2)^{\nu}}{\Gamma(\nu+1)}$ for large $\nu$) as the order $m$ increases, causing higher-order modes to vanish rapidly under moderate drift regimes. In practice, selecting a truncation order $M$ such that the contribution of omitted terms is negligible relative to the numerical integration tolerance (e.g., set to $10^{-6}$) ensures robust numerical stability. While a formal analytical error bound is reserved for future work, this exact representation serves as the basis for the numerical evaluation and performance characterization in Section IV.
@@
Figure 3: Validation of analytical CIR against Monte Carlo simulations. The analytical curves correspond to $N_{\mathrm{tx}}f_{\mathrm{3D}}^{(\mathbf{v})}(t)\,\Delta t$, i.e., the expected number of absorbed molecules within a fixed time bin $\Delta t=5\times 10^{-5}\,\mathrm{s}$. Top Row: Moderate drift speed ($|\mathbf{v}|=5\,\mu\mathrm{m/s}$). Bottom Row: Higher drift speed ($|\mathbf{v}|=10\,\mu\mathrm{m/s}$). Columns correspond to positive ($\psi=180^{\circ}$), transverse ($\psi=90^{\circ}$), and negative ($\psi=0^{\circ}$) drift directions.
@@ IV Numerical Results and Discussion
We validate the derived analytical CIR against particle-based Monte Carlo simulations and illustrate how the formulation enables efficient characterization of two key channel metrics: pulse amplitude and peak time. All simulations are conducted in MATLAB. Consistent with established MC literature [19], the default parameters are $D=80\,\mu\mathrm{m}^{2}/\mathrm{s}$, receiver radius $r=10\,\mu\mathrm{m}$, and initial distance $|\mathbf{x}_{0}|=20\,\mu\mathrm{m}$. Two representative drift magnitudes are considered: $|\mathbf{v}|=5$ and $10\,\mu\mathrm{m/s}$.
For the Monte Carlo validation in Fig. 3, the analytical CIR is discretized using a fixed time bin $\Delta t$ to match the particle-based implementation. In Figs. 4 and 5, the peak time $t_{\mathrm{peak}}$ is obtained by maximizing the continuous-time analytical CIR, and the corresponding peak amplitude is reported as $N_{\mathrm{tx}}f^{(\mathbf{v})}_{\mathrm{3D}}(t_{\mathrm{peak}})\Delta t$, using the same $\Delta t$ to ensure consistent scaling.
@@ IV-B Convergence of the Analytical Series
To evaluate (1), the angular series is truncated at order $M$. The convergence is driven by the Gaussian-like decay $e^{-\lambda^{2}t/2}$ and the factorial suppression of higher-order Legendre modes. Empirical results demonstrate that selecting $M=30$ ensures the truncation error is negligible relative to standard numerical tolerances. Although pronounced angular anisotropy (e.g., under strong drift or as $t\to 0$) may necessitate a larger $M$ to resolve fine-grained spatial features, we found that $M=30$ provides a near-optimal balance between computational efficiency and analytical precision for the performance characterization in Section IV.
@@ IV-C Characterization of Peak Metrics
Unlike computationally expensive Monte Carlo methods that require extensive averaging to suppress stochastic fluctuations, the proposed exact CIR provides a smooth objective function, enabling highly efficient, noise-free numerical maximization for peak metric extraction.
@@ V Conclusion
This letter derived an exact analytical 3-D CIR for MC systems with a FA spherical receiver under uniform drift of arbitrary direction. By formulating the problem via joint first-hitting statistics and a Girsanov measure-change argument, we obtained an exact series representation without geometric approximations or dimensional reductions. Verified via particle-based simulations, this formulation clarifies how drift orientation and receiver geometry jointly shape signal enhancement, suppression, and geometric scaling. Furthermore, as a precise and computationally tractable reference model, the proposed CIR (1) enables noise-free peak metric extraction and may facilitate advanced molecular MIMO detection and receiver design [10, 7, 1], addressing a fundamental challenge in drifted point-to-sphere channel modeling [19, 8] that has remained open since 2014.
Appendix A Measure-Change Methodological Consistency Check
This appendix provides a brief consistency check for the proposed measure-change framework (see Fig. 2). We consider the 1-D first-hitting-time problem under constant drift to demonstrate that the same measure-change argument recovers the classical IG law from the drift-free Lévy law.
In one dimension, the absorbing boundary reduces to a deterministic hitting point. Let the initial distance be $\ell>0$ and $\sigma^{2}=2D$. The no-drift first-hitting-time density follows the Lévy law [17],
$f_{\mathrm{1D}}^{(0)}(t)=\frac{\ell}{\sqrt{4\pi Dt^{3}}}\exp\!\left(-\frac{\ell^{2}}{4Dt}\right),\quad t>0.$ (7)
Under constant drift $v$ toward the boundary, the Girsanov factor at the hitting time reads: $\exp\!\big(\frac{v\ell}{2D}-\frac{v^{2}t}{4D}\big)$. Multiplying (7) by this factor yields
$f_{\mathrm{1D}}^{(v)}(t)=\frac{\ell}{\sqrt{4\pi Dt^{3}}}\exp\!\left(-\frac{(\ell-vt)^{2}}{4Dt}\right),$ (8)
thereby recovering the classical IG law [9, 17].